OpenStudy (anonymous):

solve the linear equation 12=-2(-9-3z)

OpenStudy (anonymous):

First, distribute the -2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

12=18+6z

OpenStudy (anonymous):

Good, now in order to solve for z, you have to get z completely by itself. Let's start that by subtracting 18 from both sides. What did you get?

OpenStudy (anonymous):

12=6z

OpenStudy (anonymous):

Well, you have to subtract 18 from both sides so, 12 - 18 = 18 - 18 +6z try that

OpenStudy (anonymous):

-6=6z

OpenStudy (anonymous):

Thats right good, so now to get z by itself, in this case you divide both sides by 6 What does that give you?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

Thats right! z = -1

OpenStudy (anonymous):

With these problems you can always check your answer, by substituting the value in for the variable

OpenStudy (anonymous):

thank you